30 January 2008

RHIC pp Set-Up:  It took much longer to set up the Snakes yesterday.  RF Capture was established in both rings, but without ring-to-ring synchro yet.

BPMs were timed in with fixed timing, ARTUS was set up, but we still need time for BBQ and Polarimeter Set-Up.

RHIC Chromaticity failed in Yellow.  Tunes set to nominal ramp values.  Tested store tunes in Blue.  Both rings manually decoupled. 

Both injection efficiency and lifetimes are excellent. 

Blue vertical injection damper has no effect.  Blue orbit corrections need to be scaled down by a factor 2.  Dipole trims have opposite effect than in run 6.  Blue tq polarities are OK.  Blue ARTUS kicks cause visible beam loss.  May need to reduce strength..

Ramp work on pp80 for slow factors 1 (up) and 2 (down) did not get started yesterday, so it is scheduled for immediately after the morning meeting. 

PP2PP will work behind this store. 

When beam comes back, we will finish setting up ring-to-ring synchro, and finish setting up the instrumentation BBQ and Polarimeter.

There will be work today with Rob on re-checking dampers, Yun will work on decoupling, Todd will change BPM to Gated Timing.  Late today we will begin ramp development using tune feedback. 


Scheduling Physicist: Mike Sivertz Last updated:
Email: sivertz@bnl.gov
Office Phone: 631-344-6102
Digital Pager: 7528 or 4300
Home Phone: 631-878-9020